# Structure and dynamics of electron-phonon coupled systems using neural quantum states

### v1

@@ 3.1 Ground state properties
We now turn to the calculation of two interacting electrons coupled to phonons. In cases of strong eph interactions, one can obtain a bound state of electrons, termed a bipolaron, due to the phonon-mediated attraction overcoming the Coulomb repulsion. Recent work on bipolarons in the Bond and SSH models shows that light yet mobile bipolarons exist in these models even with strong Coulomb interactions. In Fig. 6, we compare ground state energies obtained using NQS states with DQMC values reported in Ref.  97 for the Bond model. These calculations were performed on a $12\times 12$ lattice. Note that there is some cancellation of finite size errors due to attractive eph and repulsive e-e interactions. The largest NQS state here used 150 hidden neurons. For the SSH model, we performed extrapolations with respect to the number of hidden neurons to obtain the reported energies. No calculations of the binding energy have been reported for this model in the literature. However, they display a trend similar to the Bond model with the bipolarons staying bound even at large Coulomb interactions.
@@
Figure 7: Ground state energy of the Holstein-Hubbard model at half-filling on a 20 site chain with $U=4$ and $\omega_{0}=1$. Square symbols show double occupancy in the NQS state. AFQMC results taken from Ref.  29
@@
Finally, we present results for the Hubbard-Holstein model on a one-dimensional chain of 20 sites at half-filling with $U=4$. These calculations were performed in the site basis. We used a generalized Hartree Fock (GHF) state as the reference antisymmetric electronic state multiplied with an NQS eph Jastrow factor,
@@ 3.2 Dynamical properties
In Fig. 9, we show the spectral functions of the Holstein, Bond, and SSH model polaron on a 30 site chain. The spectral functions were converged with respect to the number of hidden neurons in the ansatz up to the stochastic error. For $\omega_{0}=1$ and $\lambda=2$ used here, we verified that finite size effects are negligible. In the Holstein model spectral function, the first excited state at $k=0$ is exactly $\omega_{0}$ in energy above the ground state, indicating an unbound state of a phonon well separated from the ground state polaron. We also note the appearance of a discrete state immediately above the one phonon continuum band more prominent near the band edge. In Ref.  109, this was characterized as an antibound state between the polaron and a phonon which has vanishing spectral weight at $k=0$. We do not see the nondispersive repulsive state seen in that work, which they attributed to the finite size of the lattice used in that work.
@@
We also calculated the ground and first excited states of the two-dimensional Holstein bipolaron as a function of the electronic interaction $U$. Results on a $10\times 10$ lattice are shown in Fig. 10. For moderate eph coupling, the ground state in this model evolves from a strongly bound $S_{0}$ bipolaron, with both electrons mostly on the same site, to a weakly bound $S_{1}$ bipolaron with the two electrons on neighboring sites. Our ground state energies are in good agreement with those reported in Ref.  110. In the on-site regime, the energy increases nearly linearly with $U$. Strong coupling argumentsMacridin et al. (2004) suggest the presence of two singlet excited $S_{1}$ states below the phonon and free-polaron continuua for weak to intermediate $U$. These states are bound due to the interplay of effective kinetic exchange interactions with eph coupling. One of these states has a $d$-wave symmetry, while the other has $s$-wave symmetry. Since spatial symmetry is projected in our calculations, LR-VMC on top of the $s$-wave symmetric ground state only captures excited states in this symmetry sector. We find this excited state just below the phonon continuum for small $U$. With increasing $U$ it starts mixing with the $S_{0}$ ground state and after the crossover, it becomes the ground state.
(a) Spin
(b) Charge
@@
Figure 11: Dynamical spin and charge structure factors ($\eta=0.1$) for the Hubbard-Holstein model for a half-filled 30 site chain with $\omega_{0}=5,\lambda=0.25$ and $U=4$ corresponding to the Mott insulating phase. We use a logarithmic scale. The trivial peak in the charge structure factor at $k=\omega=0$ has been removed to enhance clarity. The dashed white lines show the Bethe ansatzEssler et al. (2005) bounds for the Hubbard model with $U_{\text{eff}}=U-4\lambda=3$. For the spin structure factor, they mark the bounds of the two-spinon continuum. In the charge sector, they indicate the onset of the two-holon two-spinon (2c2s) and two-holon (2c) continuua.
In Fig. 11 we present dynamical spin and charge structure factors for the half-filled Hubbard-Holstein model on a 30 site chain. The dynamical structure factors are defined as
$\begin{split}S_{\sigma}(k,\omega)&=-\frac{1}{\pi}\text{Im}\langle\psi_{0}|S^{z}_{-k}\frac{1}{\omega-H+E_{0}+i\eta}S^{z}_{k}|\psi_{0}\rangle\\ S_{\rho}(k,\omega)&=-\frac{1}{\pi}\text{Im}\langle\psi_{0}|N_{-k}\frac{1}{\omega-H+E_{0}+i\eta}N_{k}|\psi_{0}\rangle,\end{split}$ (26)
@@
where $|\psi_{0}\rangle$ and $E_{0}$ are the ground state wave function and energy, respectively. Unlike the polaron spectral function calculations, here, we perform a single VMC optimization for the ground state at $k=0$. The LR basis functions at all $k$ points are obtained by momentum projection of the $k=0$ LR space. For the parameters used here, the system is in the Mott insulating phase with quasi-antiferromagnetic ordering. We used a $S_{z}$ and momentum projected NQS-Jastrow GHF wave function. In this phase, the spin spectrum is gapless and can be described by a two-spinon continuum, with a bulk of the spectral weight around $k=\pi$. The charge spectrum, on the other hand, is gapped and consists of two-doublon excitations. There is also a weak contribution to the structure factor due to two-doublon two-spinon excitations.Essler et al. (2005) The spectra for the Hubbard-Holstein model shown in the figure follow this expected behavior and are in good agreement with those reported in Ref.  112 obtained using the continuous time Monte Carlo method at a small but finite temperature. In the charge spectrum, we also see a feature at higher energies due to phonon excitations.
@@ Appendix B Sampling the Hamiltonian and overlap matrices
One way to mitigate this problem is to use a different sampling function. This has been recognized in various QMC excited state studies Ceperley and Bernu (1988); Filippi et al. (2009); Li and Yang (2010). Termed reweighting in Ref.  87, this method uses the following sampling approach:
@@
The statistical performance of the two approaches is shown in Fig. 13. We performed these calculations on a small Holstein chain with a truncated phonon space to allow deterministic evaluation of the spectrum in LR-VMC, which serve as reference values. We restricted the LR-VMC sampling to the same truncated Hilbert space for the sake of this comparison. While Eqs. 27 and 28 yield unbiased estimates of the $\mathbf{H}$ and $\mathbf{S}$ matrix elements, their eigenvalues are biased, as they are nonlinear functions of the matrices.Blunt et al. (2018) Using each sampling approach, we calculated averages of 100 independent calculations using various numbers of samples to compute energies of the low-lying states. As expected, there is a bias in the obtained energies for a small number of samples, which decreases systematically as we increase the number of samples in the case of reweighted sampling. On the other hand, for ground state sampling we see a persistent large bias that is nearly unchanged from around 2000 to 6000 samples and suddenly decreases for 8000 samples. This is indicative of ergodicity issues in the ground state sampling approach to LR-VMC. Here, configurations with large contributions to the excited states do not get sampled often enough if the number of samples is small, leading to a large bias, especially in the excited state energies. We also find that numerical instabilities due to linear dependencies in the basis set Blunt et al. (2018); Lee et al. (2021) are greatly reduced due to reweighted sampling.

### v2

@@ 3.1.5 Bond and SSH model bipolarons
We now turn to the calculation of two interacting electrons coupled to phonons. In cases of strong eph interactions, one can obtain a bound state of electrons, termed a bipolaron, due to the phonon-mediated attraction overcoming the Coulomb repulsion. Recent work on bipolarons in the Bond and SSH models shows that light yet mobile bipolarons exist in these models even with strong Coulomb interactions. In Fig. 8, we compare ground state energies obtained using NQS states with DQMC values reported in Ref. 101 for the Bond model. These calculations were performed on a $12\times 12$ lattice. Note that there is some cancellation of finite size errors due to attractive eph and repulsive e-e interactions. The largest NQS state here used 150 hidden neurons. For the SSH model, we performed extrapolations with respect to the number of hidden neurons to obtain the reported energies. No calculations of the binding energy have been reported for this model in the literature. However, they display a trend similar to the Bond model with the bipolarons staying bound even at large Coulomb interactions. We have performed calculations for the one-dimensional case to compare against the exact diagonalization results presented in Ref. 106 and found very close agreement, providing further validation for the accuracy of the results for the square lattice in lieu of benchmark results.
@@
Figure 9: Ground state energy of the Holstein-Hubbard model at half-filling on a 20 site chain with $U=4$ and $\omega_{0}=1$ (corresponding to the axis on the left). Square symbols show double occupancy in the NQS state (corresponding to the axis on the right). AFQMC results taken from Ref. 29
3.1.6 Holstein model with finite filling
Table 2: Ground state energy per site of the Holstein model at half-filling $\omega_{0}=1$ and $g=1.5$. Jastrow and RBM values are taken from previous VMC studies.
$N_{\text{sites}}$ JastrowOhgoe and Imada (2014) RBMNomura (2020) MLP (this work) GFMCMcKenzie et al. (1996)
4 -0.8904(5) - -0.89394(7) -0.895(1)
6 -0.8583(3) -0.868 -0.8674(2) -0.868(1)
8 -0.8424(3) -0.861 -0.8606(1) -0.861(2)
16 -0.8388(5) -0.855 -0.8535(2) -0.854(1)
We consider a Holstein model with spinless fermions that do not interact directly but only do so indirectly through the eph interaction. Previous QMC studies, including one using RBM NQS, used this model at half-filling on a one-dimensional chain as a benchmark, with Ref. 107 providing numerically exact Green’s function Monte Carlo (GFMC) ground state energies. The form of the Hamiltonian used is slightly different from the Holstein model shown in Table 1, with the eph coupling term given by
$H_{\text{eph}}=g\sum_{i}\left(c^{\dagger}_{i}c_{i}-\frac{1}{2}\right)(b^{\dagger}_{i}+b_{i}).$ (23)
We use periodic boundary conditions for systems with an odd number of fermions and antiperiodic boundary conditions otherwise, in keeping with previous studies.
Table 2 shows the ground state energy per site for different lattice sizes for $\omega_{0}=1$ and $g=1.5$. Ref. 39 used electron-phonon and conventional electron-electron Jastrow factors on top of a BCS reference wave function within VMC, whereas Ref. 66 used an RBM Jastrow on top of the same reference. The correlating factors in both these studies are non-negative and thus inherit their sign structure from the reference state. In contrast, we have used a complex NN correlator (Eq. 4) on top of a simple Hartree-Fock reference state. We impose translational invariance on this wave function by explicitly symmetrizing overlaps. This represents a more flexible ansatz with a general sign structure, but also poses a challenging problem, requiring the NN to capture the relevant physics. The difficulties in learning sign structures of many-body wave functions using NQS have been reported in the literature.Westerhout et al. (2020) Despite these challenges, we find that our method is able to obtain ground state energies in good agreement with the numerically exact GFMC results. We found that MLPs with up to $10N_{\text{sites}}$ hidden neurons were sufficient to converge the results to within the stochastic error bars of the GFMC energies. They represent a significant improvement over the conventional Jastrow results of Ref. 39 and are comparable to the converged RBM results of Ref. 66 (which were estimated from Fig. 4 of this paper).
3.1.7 Hubbard-Holstein model
@@
In this section, we present results for the Hubbard-Holstein model on a one-dimensional chain of 20 sites at half-filling with $U=4$. These calculations were performed in the site basis. We used a generalized Hartree Fock (GHF) state as the reference antisymmetric electronic state multiplied with an NQS eph Jastrow factor,
@@ 3.2.1 Convergence of the LR-VMC method
3.2.2 Spectral functions of polaron models
In Fig. 11, we show the spectral functions of the Holstein, Bond, and SSH model polaron on a 30 site chain. The spectral functions were converged with respect to the number of hidden neurons in the ansatz up to the stochastic error. For $\omega_{0}=1$ and $\lambda=2$ used here, we verified that finite size effects are negligible. In the Holstein model spectral function, the first excited state at $k=0$ is exactly $\omega_{0}$ in energy above the ground state, indicating an unbound state of a phonon well separated from the ground state polaron. We also note the appearance of a discrete state immediately above the one phonon continuum band more prominent near the band edge. In Ref. 115, this was characterized as an antibound state between the polaron and a phonon which has vanishing spectral weight at $k=0$. We do not see the nondispersive repulsive state seen in that work, which they attributed to the finite size of the lattice used in that work.
@@ 3.2.3 Low-lying states of the Holstein bipolaron
We also calculated the ground and first excited states of the two-dimensional Holstein bipolaron as a function of the electronic interaction $U$. Results on a $10\times 10$ lattice are shown in Fig. 12. For moderate eph coupling, the ground state in this model evolves from a strongly bound $S_{0}$ bipolaron, with both electrons mostly on the same site, to a weakly bound $S_{1}$ bipolaron with the two electrons on neighboring sites. Our ground state energies are in good agreement with those reported in Ref. 117. In the on-site regime, the energy increases nearly linearly with $U$. Strong coupling argumentsMacridin et al. (2004) suggest the presence of two singlet excited $S_{1}$ states below the phonon and free-polaron continuua for weak to intermediate $U$. These states are bound due to the interplay of effective kinetic exchange interactions with eph coupling. One of these states has a $d$-wave symmetry, while the other has $s$-wave symmetry. Since spatial symmetry is projected in our calculations, LR-VMC on top of the $s$-wave symmetric ground state only captures excited states in this symmetry sector. We find this excited state just below the phonon continuum for small $U$. With increasing $U$ it starts mixing with the $S_{0}$ ground state and after the crossover, it becomes the ground state.
(a) Heisenberg chain
(b) Heisenberg-Bond 2D lattice
Figure 13: (a) Convergence of the LR-VMC dynamical spin structure factor for the spin-1/2 Heisenberg model on a 32 site chain at different $k$-points with a broadening of $\eta=0.1$. $\alpha$ is the density of hidden neurons. (b) The same quantity for the Heisenberg-Bond model on a $10\times 10$ square lattice with $\omega_{0}=2$ and $g=1$. The white dashed line shows linear spin wave magnon energies with a fitted $J_{\text{eff}}=1.47J$
3.2.4 Dynamical spin structure factor of the Heisenberg and Heisenberg-Bond models
The Heisenberg model serves as a useful benchmark system for the calculation of dynamical properties since reference values can be obtained relatively easily for large systems using DMRG. This model has been used for assessing the performance of previous NQS based dynamical calculations, therefore also provides a good point of comparison. The spin-1/2 Heisenberg Hamiltonian is given by
$H=J\sum_{\langle ij\rangle}\mathbf{S}_{i}\cdot\mathbf{S}_{j},$ (27)
where $\mathbf{S}_{i}$ are spin-1/2 operators on lattice sites and the exchange interactions are limited to nearest neighbors. We set $J=1$ corresponding to an antiferromagnetic coupling. We use an NQS of the same form as Eq. 4 with the spin configuration used as the input. A quantity of interest in this model is the dynamical spin structure factor defined as
[…]
where $|\psi_{0}\rangle$ is the ground state wave function. Unlike the polaron spectral function calculations, here, we perform a single VMC optimization for the ground state at $k=0$. The LR basis functions at all $k$ points are obtained by momentum projection of the $k=0$ LR space.
We calculated $S_{\sigma}(k,\omega)$ for the spin-1/2 Heisenberg model with periodic boundary conditions on a 32 site chain at different $k$-points using LR-VMC. The results are shown in the left panel of Fig. 13. We used time-evolving block decimation to obtain the reference spectra. For LR-VMC, we used complex MLP states with an increasing number of hidden neurons with translational symmetry. Note that since this is a stoquastic Hamiltonian, it is possible to use a real MLP with the signs fixed by the Marshall sign rule. But we found the use of complex MLPs to provide faster convergence of the spectra with respect to the number of parameters. We show calculations with hidden neuron densities $\alpha=3$ and 5, demonstrating the convergence of the spectra with increasing number of parameters. We see good agreement between the LR-VMC and DMRG with some noticeable deviations especially at $k=\pi/2$ and $k=3\pi/4$. The large memory requirements of our current implementation limit the number of hidden neurons we can use in these calculations. With use of direct methods, we believe it will be possible to obtain converged results. We note that these results show an improvement over previously reported NQS calculations for this system in Ref. 69.
The power of the methods developed in this work stems from the ability to treat the coupling of localized degrees of freedom to phonons. Thus we also studied the Heisenberg-bond modelSandvik et al. (1997) which includes a phonon modulated exchange interaction given as
$H=J\sum_{\langle ij\rangle}\left(1-g(b^{\dagger}_{ij}+b_{ij})\right)\mathbf{S}_{i}\cdot\mathbf{S}_{j}+\omega_{0}\sum_{ij}b^{\dagger}_{ij}b_{ij},$ (29)
where $b_{ij}$ are phonon destruction operators residing on the lattice bonds. We again set $J=1$. The dynamical spin structure factor for this model on a $10\times 10$ lattice is shown in the right panel of Fig. 13. We used antiperiodic boundary conditions for this lattice. The results were obtained using LR-VMC with a complex MLP state with a density of hidden neurons $\alpha=2$. We see a magnon band similar to the bare Heisenberg model. Fitting the linear spin wave theory result to the peaks in the LR-VMC magnon dispersion, we find $J_{\text{eff}}=1.47J$, indicating a renormalized exchange interaction. We note that this scaling of $J$ also partially reflects the underestimation of magnon energies by the linear spin wave theory. There is an additional structure around the optical phonon frequency $\omega_{0}$ on top of the magnon spectrum due to the spin-phonon interaction. Thus for this set of parameters, the ground state is an antiferromagnet with renormalized exchange interactions due to the phonon coupling. We emphasize that the approach outlined here gives access to spectral properties of systems in two and higher dimensions coupled to phonons which are challenging for other numerical approaches like DMRG. We leave a detailed study of different phases of this system over the full parameter space to future work.
(a) Hubbard chain
(b) Hubbard-Holstein chain
@@ 3.2.4 Dynamical spin structure factor of the Heisenberg and Heisenberg-Bond models
Figure 14: (a) Convergence of the LR-VMC dynamical charge structure factor for the half-filled Hubbard model ($U=8$) on a 14 site chain at different $k$-points with a broadening of $\eta=0.2$. $\alpha$ is the density of hidden neurons. (b) Spin and charge structure factors for the Hubbard-Holstein model for a half-filled 30 site chain with $\omega_{0}=5,\lambda=0.25$ and $U=4$. The trivial peak in the charge structure factor at $k=\omega=0$ has been removed to enhance clarity. The dashed white lines show the Bethe ansatzEssler et al. (2005) bounds for the Hubbard model with $U_{\text{eff}}=U-4\lambda=3$. For the spin structure factor, they mark the bounds of the two-spinon continuum. In the charge sector, they indicate the onset of the two-holon two-spinon (2c2s) and two-holon (2c) continuua.
3.2.5 Dynamical properties of the Hubbard-Holstein model
The Hubbard-Holstein model allows us to study the effect of eph coupling on charge dynamics in addition to the spin dynamics. Besides the dynamical spin structure factor, defined in the same way as for the Heisenberg model (Eq. 28), we consider the charge structure factor defined as
$S_{\rho}(k,\omega)=-\frac{1}{\pi}\text{Im}\langle\psi_{0}|N_{-k}\frac{1}{\omega-H+E_{0}+i\eta}N_{k}|\psi_{0}\rangle,$ (30)
where $|\psi_{0}\rangle$ and $E_{0}$ are the ground state wave function and energy, respectively. We first perform benchmark calculation on the purely electronic Hubbard model without eph coupling, where it is easier to obtain near-exact reference results for systems of nontrivial sizes. The left panel of Fig. 14 shows the convergence of the charge structure factor for the half-filled 14 site Hubbard model with $U=8$ using LR-VMC. We compare our results to the reference ED values presented in Ref. 96. We confirmed that we obtain identical results for ED at $k=\pi$ using our code. We used a $S_{z}$ and momentum projected NQS-Jastrow GHF wave function. This represents a challenging regime for the LR-VMC method as it requires a good description of higher energy charge excitations. Despite this we see good agreement with the ED results for $\alpha=7$ which uniformly improves upon the $\alpha=4$ calculation.
@@ 3.2.5 Dynamical properties of the Hubbard-Holstein model
In the right panel of Fig. 14 we present dynamical spin and charge structure factors for the half-filled Hubbard-Holstein model on a 30 site chain. For the electronic part, we used the same wave function as the Hubbard model calculation, with phonons correlated in the NQS Jastrow factor. For the Hamiltonian parameters used here, the system is in the Mott insulating phase with quasi-antiferromagnetic ordering. In this phase, the spin spectrum is gapless and can be described by a two-spinon continuum, with a bulk of the spectral weight around $k=\pi$. The charge spectrum, on the other hand, is gapped and consists of two-doublon excitations. There is also a weak contribution to the structure factor due to two-doublon two-spinon excitations.Essler et al. (2005) The spectra for the Hubbard-Holstein model shown in the figure follow this expected behavior and are in good agreement with those reported in Ref. 120 obtained using the continuous time Monte Carlo method at a small but finite temperature. In the charge spectrum, we also see a feature at higher energies due to phonon excitations.
@@ Appendix B Sampling the Hamiltonian and overlap matrices
One way to mitigate this problem is to use a different sampling function. This has been recognized in various QMC excited state studies Ceperley and Bernu (1988); Filippi et al. (2009); Li and Yang (2010). Termed reweighting in Ref. 91, this method uses the following sampling approach:
@@
The statistical performance of the two approaches is shown in Fig. 16. We performed these calculations on a small Holstein chain with a truncated phonon space to allow deterministic evaluation of the spectrum in LR-VMC, which serve as reference values. We restricted the LR-VMC sampling to the same truncated Hilbert space for the sake of this comparison. While Eqs. 31 and 32 yield unbiased estimates of the $\mathbf{H}$ and $\mathbf{S}$ matrix elements, their eigenvalues are biased, as they are nonlinear functions of the matrices.Blunt et al. (2018) Using each sampling approach, we calculated averages of 100 independent calculations using various numbers of samples to compute energies of the low-lying states. As expected, there is a bias in the obtained energies for a small number of samples, which decreases systematically as we increase the number of samples in the case of reweighted sampling. On the other hand, for ground state sampling we see a persistent large bias that is nearly unchanged from around 2000 to 6000 samples and suddenly decreases for 8000 samples. This is indicative of ergodicity issues in the ground state sampling approach to LR-VMC. Here, configurations with large contributions to the excited states do not get sampled often enough if the number of samples is small, leading to a large bias, especially in the excited state energies. We also find that numerical instabilities due to linear dependencies in the basis set Blunt et al. (2018); Lee et al. (2021) are greatly reduced due to reweighted sampling.